Express v=(2,-5,3) in R3 as a linear combination of the vectors U1 =(1,-3,2) , U2=(2,-4,-1), U3 =(1,-5,7)


1
Expert's answer
2021-03-05T01:26:58-0500

Rewriting the equation in the form,

"\begin{bmatrix}
 2 \\
 -5\\
 3
\end{bmatrix}= x\begin{bmatrix}
 1 \\
 -3\\
 2
\end{bmatrix}+y\begin{bmatrix}
 2 \\
 -4\\
 -1
\end{bmatrix}+z\begin{bmatrix}
 1 \\
 -5\\
 7
\end{bmatrix}\\
\begin{bmatrix}
 x + 2y + z \\
 -3x -4y -5z \\
 2x -y +7z
\end{bmatrix}=\begin{bmatrix}
 2 \\
 -5\\
 3
\end{bmatrix}\\
\text{reducing the matrix from the back yields}\\.
\begin{bmatrix}
 x + 2y + z=2 \\
 -3x -4y -5z=-5 \\
 2x -y +7z=3
\end{bmatrix}\\
-2R_1+R_3 \implies R_3\\
3R_1+R_2 \implies R_2\\
\begin{bmatrix}
 x + 2y + z=2 \\
 2y -2z=1 \\
 -5y +5z=-1
\end{bmatrix}\\
\frac{5}{2}R_2+R_3 \implies R_3\\
\\
\begin{bmatrix}
 x + 2y + z=2 \\
 2y -2z=1 \\
 0=\frac{3}{2}
\end{bmatrix}\\
\text{The system has no solution.}\\
\text{Thus it is impossible to write the vector v as a linear combination of}\\u_1, u_2, u_3"